Tug the macaroni penguin throws a frozen 12.00 g krill toward a
vertical wall. The frozen krill strikes the vertical wall traveling with a
speed of 8.00 m/s. The frozen krill rebounds off of the wall horizontally
and the force that the wall exerts on the krill during the time of contact
is shown in the graph (rightwards is the positive direction).
a. Calculate the impulse imparted on the frozen krill.
b. Find the equivalent average force exerted on the frozen krill.
c. Find the rebound velocity of the frozen krill.

ca FW. 0.48 0,24 tel. -424+
Answer #1

a) Here net impulse j = (0.48) X004 +60.24) X0.2 =0,192 - 0048 = 0.144 NS b) If equivalent average fouce is F then - Ext = J